Please help me solve this word problem:
Stu has twice as much money invested in stocks as in bonds. Stocks earn 8% interest per year and bonds 2% per year. If Stu earned a total of $400 dollars from his stocks and bonds last year, how much money did he have invested in stocks?
----- Thank you =)

Answer by checkley77(12844)

.08*2x+.02x=400
.16x+.02x=400
.18x=400
x=400/.18
x=2,222.22 invested @ 2% in bonds.
2*2,222.22=4,444.44 invested @ 8% in stocks.
Proof:
.08*4,444.44+.02*2,222.22=400
355.55+44.44=400
399.99~400